OpenStudy (anonymous):

Use an appropriate change of base formula to convert the following expression to one with the indicated base? log(x^6 + 5) to base e?

OpenStudy (anonymous):

\[\frac{\ln(x^6+5)}{\ln(10)}\] assuming log meant log base ten

OpenStudy (anonymous):

yess

OpenStudy (anonymous):

do you do anything after that?

OpenStudy (anonymous):

nothing else you can do

OpenStudy (anonymous):

change of base says \[\log_a(x)=\frac{\log_b(x)}{\log_b(a)}\] which is another way of saying that if you want to solve \[a^y=x\] for y you can say \[y=\log_a(x)\] or you can say \[y=\frac{ln(y)}{ln(a)}\]which has the advantage that you can actually use a calculator and compute your answer

OpenStudy (anonymous):

In general, let a, b, and x be positive real numbers such that \[a,b \neq1\]Then \[\log_{a} x\]can be converted to the base b by the formula\[\frac{\log_{b} x}{\log_{b} a}\]

OpenStudy (anonymous):

i see it now thanksss

OpenStudy (anonymous):

Use an appropriate change of base formula to convert the following expression to one with the indicated base.log(x^8 + 7)to base

OpenStudy (anonymous):

can u help me